Problem 4 and 5-4 Future Value and Number of Annuity Payments Your client has been given a trust fund valued at $1.61 million please dont round decimals until the final answer
Answer #1

FV = PV x (1 + r)n = 1,160,000 x (1 + 4%)20 = $ 2,541,703

This will be the size of kitty available for withdrawal at the time of retirement.

Hence, at the time of retirement, PV = $ 2,541,703

Monthly withdrawals = PMT = -19,500

Rate = interest rate / month = 4% / 12 = 0.333%

Fv = 0

Hence, Number of months this periodical withdrawal can continue = NPER (Rate, PMT, PV, FV) = NPER (0.333%, -19500, 2541703, 0) = 171.22 months

Hence, please enter 171.22 in the answer box.